OpenStudy (anonymous):

prove that the series 0.6+0.06+0.006+... converges and find it sum. i found 0.1 for the ratio and 0.667 for the sums.It's correct?

OpenStudy (anonymous):

this is the same as \(\overline{.6}\) which is also \(\frac{2}{3}\)

OpenStudy (anonymous):

if you want to add a series, then use \(\frac{a}{1-r}\) with \(a=\frac{6}{10}, r=\frac{1}{10}\)

OpenStudy (anonymous):

Thanks